# Assessment of DKIST/VTF Capabilities for the Detection of Local Acoustic Source Wavefronts

### v1

@@ 5 Spectral signatures
The first step in quantifying the spectral sensitivity to local wavefront perturbations is the synthesis of the emergent Stokes $I$ intensity from the MURaM simulation. We synthesize three photospheric lines: Fe I 5250.2 Å, Fe I 5250.6 Å, and Fe I 6302.5 Å, over wavelength ranges 5249.90–5250.45 Å, 5250.45–5251.00 Å, and 6302.20–6302.90 Å with 0.005 Å sampling. We consider only vertical rays and restrict the vertical extent of the atmosphere to the unshaded region in Figure 2 ($0.30\leq z\leq\,1.28$ Mm). The latter restriction reduces the computational cost of the transfer calculation while still capturing the full range of heights that contribute to absorption in the lines of interest. Using the temperature, pressure, and line-of-sight velocity profiles with height in a small region of interest (see Section 5.2 below), we synthesize emergent spectra for each atmospheric column. We employ the SNAPI code (Milić and van Noort, 2018), solving the polarized radiative transfer equation assuming Local Thermodynamic Equilibrium (LTE). Thus, we obtain $I(\lambda,t)$ for each atmospheric column, at each horizontal location in the simulation.
@@ 5.1 Response
The sensitivity of the emergent intensity to atmospheric perturbations is quantified through the spectral response function. Of particular interest for wavefront detection is the temperature response function (${\rm RF}_{T}$), which measures the linear changes in the emergent intensity at a given frequency produced by localized perturbations in $T$ as a function of perturbation height $z$. SNAPI computes the derivatives of the specific intensity with respect to changes in $T$ by evaluating variations in the formal solution that result from small perturbations to the atmospheric column at each height. The result is a two–dimensional kernel,
@@
The response functions for the wavelength ranges we consider peak between $z\approx 0\,$km and 500500\,km. Because the acoustic wavefronts in the simulation are most apparent as small amplitude fluctuations in the overshoot region, which lies within this height range, these response functions provide a natural bridge between the physical perturbations induced by the wavefront passage and its observational signatures.
@@ 5.2 Perturbation definition and wavefront mask
With the goal of determining where in the spectrum the acoustic wavefront will leave its largest signature, we first identify the location of the wavefront as a function of time in the simulation. We then find the wavelengths for which the signal from those locations in the atmosphere is largest compared to other background contributions (Section 5.3). This will motivate observation strategies that either fix on the most sensitive wavelengths or amplify the wave signature by tracking the wave in spectral space (Section 6).
@@
To determine the locations at which the wavefront is most clearly defined against the background, we manually locate its position at each height as a function of time in the $v^{(3)}$ (velocity difference) images. We measure its angular extent and, by fitting circular arcs to its shape, locate the center of its expansion. Using these measurements, we define the region of wavefront propagation (the wedge in Figure 3A) and divide it into concentric annular segments with fixed radial widths of 3 pixels (48 km). Three pixels is the approximate radial extent of the wavefront in the direction of propagation in the simulations.
@@
Figure 3: Construction of azimuthally averaged acoustic mask from the MURaM simulation. (A) Temporal velocity difference $v^{(3)}$ snapshot (left) showing the acoustic wavefront in the MURaM photosphere with concentric annuli around fitted center (cyan). Annulus selected for analysis indicated with dark, bolded boundaries. (B) Standard score Z(z,t)Z(z,\,t) map for the region of interest (right), revealing a clear, height-dependent acoustic propagation pattern. Dashed black lines indicate the most wave-sensitive height range while the solid red line outlines the region isolated by the mask described in Section 5.2.
Focusing on temperature fluctuations (with similar analysis also done for Doppler velocity and pressure perturbations), we compute, at each height $z$ and time $t$, the average temperature in each annulus in the wavefront region, $\bar{T}(r,z,t)$, where $r$ labels the distance of the annulus from the center of expansion. Using this, we define perturbations as
$\delta T(r,z,t)=\bar{T}(r,z,t)-\bar{T}_{\mathrm{bg}}(r,z),$ (2)
where $\bar{T}_{\mathrm{bg}}$ is the mean temperature in each annulus at each height averaged over a wave-free time interval (simulation time period 150415601504-1560\,s, ending approximately 3030\,s before the wavefront is visible at any $r$ in the photosphere). It is important to note that the temperature perturbation $\delta T(r,z,\,t)$ is defined relative to the mean quiet background state, and thus does not explicitly separate the wavefront contributions from others such as granulation. As the background thermodynamic structure evolves in time, $\delta T(r,z,\,t)$ does not represent a purely wave-isolated perturbation but the total deviation from the mean quiet reference atmosphere when the wave is present.
To select the annulus with the strongest wave signal for further analysis, we compute the $v^{(3)}$ standard score ($Z$-score) for each annulus, and select the annulus with the strongest integrated $Z$-score over height and time. This annulus is highlighted with bolded boundaries in Figure 3A. While subsequent analysis focuses primarily on temperature perturbations and their influence on the emergent spectra, we defined this annular region of interest using $v^{(3)}$ because the acoustic wavefront shows the strongest signal in this variable. As illustrated by Figure 4, which plots the average $v^{(3)}$ and $\delta T^{(3)}$ in the annulus of interest as a function of time and height, and as expected for a non-oscillatory acoustic pulse, there is little phase difference between $v^{(3)}$ and $T^{(3)}$ (and thus the wave-induced velocity and temperature perturbations) as the wavefront propagates.
@@
Figure 4: One-dimensional upward wave propagation constructed from azimuthally averaged perturbations in the annulus of interest. The temporally differenced velocity $v^{(3)}$ (solid black line) as well as differenced temperature $\delta T^{(3)}$ (dashed black line) are shown as functions of time and vertically offset by height, revealing the coherent upward progression of the acoustic wavefront through the atmosphere. The shaded red area indicates the wave mask $M_{\mathrm{wave}}$ of Section 5.2 used to isolate the wave-dominated portion of the signal.
@@
With the annulus of interest defined, a time-height map of the $Z$-score for that annulus can be used to isolate the wave perturbation. The upward propagating wavefront in this map (Figure 3B) is clearly visible in the convective overshoot region, corresponding to the region where $|Z|>3$ between $0.0\leq z\leq 0.42$ (bounded by horizontal fiducial linesin the plot). Thus, a mask $M_{\mathrm{wave}}(z,t)$ that selects the time steps surrounding the $Z$-score peak at each height can be used to isolate the wavefront. We take one time-step either side of that peak ($\pm 2\,$s) to define when the wave is present at each height because this approximates the time it takes for the acoustic perturbation to cross three radial grid cells ($\sim 48\,$km), which is about the thickness of the wavefront in the direction of propagation. This time-height mask is over-plotted with a red outline in Figure 3B and superimposed as a shaded region in Figure 4. Although the masked region looks like a trough in the difference variables, it physically corresponds to a region of positive (upward) velocity and positive temperature perturbation. With this mask, the wave temperature perturbations can be identified as
@@
This is the characteristic temperature perturbation at the wave location as a function of height and time.
@@ 5.3 Sensitivity Measure
For small amplitude temperature perturbations, the change in emergent intensity at each horizontal location in the simulation at a given wavelength $\lambda$ is approximately
$\Delta I(\lambda,t)=\int\mathrm{RF}_{T}(\lambda,z)\,\delta T(z,t)\,dz\ ,$ (4)
where contributions from velocity and pressure fluctuations are not included because their contributions to $\Delta I$ are negligible for the photospheric lines we are considering. With Equation 4, we can define a wavelength-dependent sensitivity measure $\mathrm{SNR}(\lambda,t)$ to assess the wavelengths at which the wavefront signal exceeds most that due to typical background fluctuations. The wave-related contributions to the intensity changes $\Delta I(\lambda,t)$ are
@@
$\mathrm{SNR}(\lambda,t)=\frac{\Delta I_{\mathrm{wave}}(\lambda,t)}{\Delta I_{\mathrm{std}}(\lambda,t)}\ ,$ (7)
@@
$\lambda^{\ast}(t)=\underset{\lambda}{\arg\max}\;\mathrm{SNR}(\lambda,t)\ .$ (8)
As with the response function (Equation 1), the wavefront sensitivity measure (Equation 7), and thus $\lambda^{\ast}(t)$, is computed independently for each pixel in the annulus of interest. The average $\lambda^{\ast}(t)$, over that annulus, is plotted as function of time in Figure 5, and represents the wavelength at which the characteristic signal from the wavefront has the largest amplitude compared to variations at that wavelength due to fluctuations at all heights.
@@
We define a second quantity of interest $\lambda^{**}$, the spatial average (over the annulus of interest) of the wavelength at which the sensitivity measure is highest at any time. For each spatial pixel within the annulus of interest, we find the wavelength at which the wave sensitivity reaches its global maximum when considering all wavelengths and all times within the wave mask. We report a single $\lambda^{\ast\ast}$ for each spectral line by spatially averaging this quantity, yielding the wavelength at which the wavefront is, on average, most detectable. The distribution of wavelength values contributing to this mean (the individual pixel values) is shown in Figure 6. The values found for $\lambda^{**}$ and their standard deviations are $\lambda^{**}=5250.148\,$Å , 5250.583 Å , 6302.425 Å , and $\sigma_{\lambda^{**}}=0.0031$ Å, $0.0028$ Å, $0.0046$ Å  for the Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5 lines respectively. For all three lines, the maximum sensitivity occurs at $t=1642$ s. This time corresponds to a wavefront height between $z=0.128$ and $0.176$ Mm above the photosphere (above the mean $\tau=1$ height in the simulation).
Figure 6: The average, over the annulus of interest, of the temperature response function $\mathrm{RF}_{T}$ for each atmosphere making up the annulus of interest at t=1640t=1640\,s, across the Fe I 5250.2, 5250.6, and 6302.5 spectral lines (background) The probability density function of $\lambda^{\ast\ast}$ for each pixel in the annulus of interest is over-plotted in red, with $\lambda^{\ast\ast}$ indicated with a black vertical line. The insets zoom into the $\lambda^{\ast\ast}$ region and indicate the VTF spectral element extent with dotted vertical lines.
@@ 5.4 Wavefront signatures
As the wave propagates upward, the wavelength of maximum sensitivity $\lambda^{*}(t)$ drifts to longer wavelengths with time for all the lines examined (Figure 5). For all of the lines, the maximum spectral sensitivity occurs on the blue (shorter wavelength) side of the line. This is due to the systematic Doppler distortion of the lines caused by the granular upflow underlying the region of wavefront propagation. As shown in Figure 6, the Doppler distortion of the line decreases the slope of the height of the response function peak with wavelength. With the decreased slope, a narrower range of heights contributes to the response at any given wavelength on the short wavelength side of the spectral line than on its right, increasing the sensitivity measure that we have defined. In general, line profiles with less steep wings are preferred when localizing observations in height.
Because maximum spectral sensitivity occurs on the short wavelength side of the line, the wavelength of maximum sensitivity shifts to longer wavelengths (towards the line core) as the wavefront propagates upward in the atmosphere; the location of the wavefront perturbation, captured by the mask, moves higher in the atmosphere with time (Figure 3B) so that wavelength of maximum sensitivity moves toward the line core. The rate of this systematic shift does not simply reflect the wavefront propagation speed. Instead, $\lambda^{\ast}(t)$ is determined by the atmosphere through which the wave is traveling (defining the response function), the vertical propagation speed of the wavefront, and the spectral slope of the line wing. It can be used in the development of a robust observing strategy (Section 6), either by fitting $\lambda^{\ast}(t)$ for each spectral line using the simulation or, more satisfactorily, by determining from first principles how $\lambda^{\ast}$ evolves given the spectral line and the atmosphere of interest. The latter is the focus of future efforts. Here we measure the slopes and find them to be $9.223\times 10^{-04}$, $9.845\times 10^{-04}$, and $1.621\times 10^{-03}\,$Å/s  for Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5 respectively.

### v2

@@ 5 Spectral signatures
The first step in quantifying the spectral sensitivity to local wavefront perturbations is the synthesis of the emergent Stokes $I$ intensity from the MURaM simulation. Given that the heights of interest (see Figure 2) lie in the region between the base of the photosphere and the temperature minimum (the region from about 0-0.5 Mm above mean $\log\tau=0$), we closely examine three photospheric lines, synthesizing Fe I 5250.2 Å, Fe I 5250.6 Å, and Fe I 6302.5 Å, over wavelength ranges 5249.90–5250.45 Å, 5250.45–5251.00 Å, and 6302.20–6302.90 Å with 0.005 Å sampling. All three lines are commonly chosen for their magnetic sensitivity (not central to our study) and have been observed with imaging spectropolarimeters such as CRISP (Scharmer et al., 2008) and TuMag (del Toro Iniesta et al., 2025). In our synthesis, we consider only vertical rays and restrict the vertical extent of the atmosphere to the unshaded region in Figure 2 ($0.30\leq z\leq\,1.28$ Mm). The latter restriction reduces the computational cost of the transfer calculation while still capturing the full range of heights that contribute to absorption in the lines of interest. Using the temperature, pressure, and line-of-sight velocity profiles in a small region of interest (see Section 5.2 below), we synthesize emergent spectra for each atmospheric column. We employ the SNAPI code (Milić and van Noort, 2018), solving the polarized radiative transfer equation assuming Local Thermodynamic Equilibrium (LTE). Thus, we obtain $I(\lambda,t)$ for each atmospheric column at each horizontal location in the region of interest.
@@ 5.1 Response
The sensitivity of the emergent intensity to atmospheric perturbations is quantified through the spectral response function. Of particular interest for wavefront detection is the temperature response function (${\rm RF}_{T}$), which measures the linear changes in the emergent intensity at a given wavelength produced by localized perturbations in $T$ as a function of perturbation height $z$. SNAPI computes the derivatives of the specific intensity with respect to changes in $T$ by evaluating variations in the formal solution that result from small perturbations to the atmospheric column at each height. The result is a two–dimensional kernel,
@@
The response functions for the wavelength ranges we consider peak between $z\approx 0\,$Mm and 0.50.5\,Mm. Because the acoustic wavefronts in the simulation are most apparent as small amplitude fluctuations in the overshoot region, which lies within this height range, these response functions provide a natural bridge between the physical perturbations induced by the wavefront passage and their observational signatures.
@@ 5.2 Perturbation definition and wavefront mask
With the goal of determining where in the spectrum the acoustic wavefront will produce the largest observable signature, we first locate the wavefront in the simulation by manually tracing its position in the $v^{(3)}$ (velocity difference) images at the photosphere ($z=0$ Mm). We measure its angular extent and locate the center of its expansion by fitting semicircular arcs to its shape. Using these measurements, we define the region of wavefront propagation (the wedge in Figure 3A) and divide it into concentric annular segments with fixed radial widths of 3 pixels (48 km), which is the approximate extent of the wavefront in the direction of propagation in the simulations.
For each annulus, we compute the $v^{(3)}$ standard score ($Z$-score) at each height and time and define the wave as present when $|Z|>3$. The annulus with the strongest integrated $Z$-score over height and time is selected here for further analysis and is highlighted with bold boundaries in Figure 3A. A time-height map of the $Z$-score for this annulus is shown in Figure 3B, where the upward-propagating wavefront is clearly visible in the convective overshoot region, the region between $0.0\leq z\leq 0.42$ and bounded by horizontal fiducial lines in the figure. As an aside, the figure also shows evidence of a weaker possible downward propagating feature that may be relevant to observations of downward propagating waves in the solar atmosphere (e.g., Guevara Gómez et al., 2022; Chaturmutha et al., 2024, and references therein) and perhaps related to shock dissipation (Bahauddin et al., 2024), but it has not yet been analyzed in detail and is not the subject of this paper.
@@
Figure 3: Construction of azimuthally averaged acoustic mask from the MURaM simulation. (A) Temporal velocity difference $v^{(3)}$ snapshot (left) showing the acoustic wavefront in the MURaM photosphere with concentric annuli around fitted center (cyan). Annulus selected for analysis indicated with dark, bolded boundaries. (B) Standard score Z(z,t)Z(z,\,t) map for the region of interest (right), revealing a clear, height-dependent acoustic propagation pattern. The wavefront becomes clearly detectable in the selected annulus at 1632 s and remains visible for 50 s. The simulation interval 1584–1682 s is chosen for spectral synthesis to include a pre-event interval of comparable duration with no visible wavefront signal. Dashed black lines indicate the most wave-sensitive height range while the solid red line outlines the region isolated by the mask described in Section 5.2.
@@
Figure 4: One-dimensional upward wave propagation constructed from azimuthally averaged perturbations in the annulus of interest. The temporally differenced velocity $v^{(3)}$ (solid black line) and differenced temperature perturbation $\delta T^{(3)}$ (dashed black line) are shown as functions of time and vertically offset by height, revealing the coherent upward progression of the acoustic wavefront through the atmosphere. The shaded red area indicates the wave mask $M_{\mathrm{wave}}$ of Section 5.2 used to isolate the wave-dominated portion of the signal.
Within the selected annulus, we compute the azimuthally averaged temperature $\bar{T}(z,t)$ at each height $z$ and time $t$ and define the perturbation relative to a wave-free reference as
$\delta T(z,t)=\bar{T}(z,t)-\bar{T}_{\mathrm{bg}}(z),$ (2)
where $\bar{T}_{\mathrm{bg}}$ is the annular mean temperature at each height averaged over a wave-free time interval (simulation time period 150415601504-1560\,s, ending approximately 3030\,s before the wavefront is visible in any annulus in the photosphere). Analogous calculations are performed for Doppler velocity and pressure.
@@
To isolate the temperature perturbation associated with the wavefront, we define a mask $M_{\mathrm{wave}}(z,t)$ by selecting the time steps surrounding the $Z$-score peak ($\pm 2\,$s) at each height. This approximates the time it takes for the acoustic perturbation to cross 3 grid cells (4848\,km), which is about the thickness of the wavefront in the direction of propagation. The mask is shown as a red outline in Figure 3B and is superimposed as a shaded region in Figure 4. Although the masked region looks like a trough in the temporally differenced variables, it physically corresponds to a region of positive (upward) velocity and positive temperature perturbation.
Using this mask, we thus define the characteristic temperature perturbation at the wave location as a function of height and time as
@@
We note that the temperature perturbation $\delta T_{\mathrm{wave}}(z,\,t)$ is defined relative to the mean quiet background state, and thus does not explicitly separate the wavefront contributions from others such as granulation. As the background thermodynamic structure evolves in time, $\delta T_{\mathrm{wave}}(z,\,t)$ does not represent a purely wave-isolated perturbation but the total deviation from the mean quiet reference atmosphere when the wave is present.
Subsequent analysis focuses on the temperature perturbations caused by the wavefront passage because they contribute most strongly to changes in the spectrum. However, we define the location of the wavefront using $v^{(3)}$ because it is most easily identified in the simulations using this measure, and accurately locating the wavefront is critical to our assessment of its observability. For this excitation event, the wave-induced temperature and Doppler velocity perturbations are in phase at the wavefront location (Figure 4). Thus, the location of the wavefront based on $v^{(3)}$ is a consistent proxy for the location of the wave-induced temperature perturbations. The temperature-velocity phase relation found here is somewhat surprising in light of previous studies of acoustic waves on the Sun, which report a phase difference between velocity and temperature or intensity (e.g., Marmolino and Severino, 1990; Straus and Bonaccini, 1997; Rast and Bogdan, 1998; Straus et al., 1999; Skartlien and Rast, 2000; Severino et al., 2001; Vesa et al., 2023). Those previous results pertain to continuously driven modal oscillations and atmospheric wave fields for which the phase relation can be different from that of the local response to a discrete impulsive source.
@@ 5.3 Sensitivity Measure
For small amplitude perturbations, the change in emergent intensity at each height in the simulation at a given wavelength $\lambda$ is approximately
$\Delta I(\lambda,t)=\int\left[\mathrm{RF}_{T}(\lambda,z)\,\delta T(z,t)+\mathrm{RF}_{v}(\lambda,z)\,\delta v(z,t)+\mathrm{RF}_{p}(\lambda,z)\,\delta p(z,t)\right]\,dz\ .$ (4)
We exclude contributions from velocity and pressure fluctuations in our analysis because the temperature contribution to $\Delta I$ dominates for the spectral lines and wavefront event we consider here (see Figure 5).
Figure 5: Ratios of the amplitudes of the maximum intensity perturbations due to velocity and temperature fluctuations, i.e., $R(\lambda)=\mathrm{max}_{t}[\Delta I_{v}(\lambda,t)]/\mathrm{max}_{t}[\Delta I_{T}(\lambda,t)]$, for the specific acoustic event considered here.
With Equation 4, we can define a wavelength-dependent sensitivity measure $\mathrm{SNR}(\lambda,t)$ to assess at which wavelengths the wavefront signal most exceeds that due to typical background fluctuations. The wave-related contributions to the intensity changes $\Delta I(\lambda,t)$ are
@@
$\mathcal{S}(\lambda,t)=\frac{\Delta I_{\mathrm{wave}}(\lambda,t)}{\Delta I_{\mathrm{std}}(\lambda,t)}\ ,$ (7)
@@
$\lambda^{\ast}(t)=\underset{\lambda}{\arg\max}\;\mathcal{S}(\lambda,t)\ .$ (8)
As with the response function (Equation 1), the wavefront sensitivity measure $\mathcal{S}(\lambda,t)$, and thus $\lambda^{\ast}(t)$, is computed independently for each pixel in the annulus of interest. The average $\lambda^{\ast}(t)$ over that annulus is plotted as function of time in Figure 6, and represents the wavelength at which the characteristic signal from the wavefront has the largest amplitude compared to variations at that wavelength due to fluctuations at all heights.
@@
We define a second quantity of interest, $\lambda^{**}$, as the spatial average (over the annulus of interest) of the wavelength at which the sensitivity measure is highest at any time. For each pixel within the annulus of interest, we find the wavelength at which the wave sensitivity is maximum when considering all wavelengths and all times within the wave mask. We report a single $\lambda^{\ast\ast}$ for each spectral line by spatially averaging this quantity, yielding the wavelength at which the wavefront is, on average, most detectable.
The distribution of wavelength values contributing to this mean (the individual pixel values) is shown in Figure 7. The values found for $\lambda^{**}$ and their standard deviations are $\lambda^{**}=5250.148\,$Å , 5250.5835250.583\,Å, 6302.4256302.425\,Å , and $\sigma_{\lambda^{**}}=0.0031$ Å, $0.0028$ Å, $0.0046$ Å  for the Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5 lines, respectively. The time at which the sensitivity measure $\mathcal{S}(\lambda,t)$ reaches its maximum (i.e., when we measure the largest change in the emergent intensity as the wavefront passes) is $t=1642$ s, corresponding to a wavefront height between $z=0.128$ and $0.176$ Mm above the photosphere (above the mean $\tau=1$ height in the simulation). Because all three lines considered belong to the same species and exhibit comparable line strengths that lead to relatively similar sensitivities, perturbations to their intensities evolve similarly as the wavefront propagates through the atmosphere. As a result, the sensitivity measure peaks at the same time for all three lines.
Figure 7: The average, over the annulus of interest, of the temperature response function $\mathrm{RF}_{T}$ (normalized by continuum intensity $I_{c}$) for each atmosphere making up the annulus of interest at t=1640t=1640\,s, across the Fe I 5250.2, 5250.6, and 6302.5 spectral lines (background) The probability density function of $\lambda^{\ast\ast}$ for each pixel in the annulus of interest is over-plotted in red, with $\lambda^{\ast\ast}$ indicated with a black vertical line. The insets zoom into the $\lambda^{\ast\ast}$ region and indicate the VTF spectral element extent with dotted vertical lines.
@@ 5.4 Wavefront signatures
For all lines examined, maximum spectral sensitivity occurs on the blue (shorter wavelength) side of the line, due to the systematic Doppler distortion caused by the granular upflow underlying the region of wavefront propagation. As shown in Figure 7, this distortion reduces the slope of the response function, resulting in a relatively narrow range of heights which contribute to the emergent intensity. This, in turn, increases the sensitivity measure $\mathcal{S}(\lambda,t)$ given in Equation 7. Thus, line profiles with less steep wings are generally preferred because they localize the response function in height and reduce contamination from noise-dominated regions of the atmosphere.
As shown in Figure 6, maximum spectral sensitivity shifts towards the line core as the wavefront propagates upward in the atmosphere. The rate of this systematic shift does not simply reflect the wavefront propagation speed. Instead, $\lambda^{\ast}(t)$ is determined by the atmosphere through which the wave is traveling (defining the response function), the vertical propagation speed of the wavefront, and the spectral slope of the line wing. It can be used in the development of a robust observing strategy (Section 6), either by fitting $\lambda^{\ast}(t)$ for each spectral line using the simulation or, more satisfactorily, by determining from first principles how $\lambda^{\ast}$ evolves given the spectral line and the atmosphere of interest. The latter is the focus of future efforts. Here we measure the slopes and find them to be $9.223\times 10^{-4}$, $9.845\times 10^{-4}$, and $1.621\times 10^{-3}\,$Å s-1  for Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5 respectively.